OpenStudy (anonymous):

A baseball is hit inside a baseball diamond with a length and width of 90 feet each. What is the probability that the ball will bounce on the pitcher’s mound, if the diameter of the mound is 18 feet? Assume that the ball is equally likely to bounce anywhere in the infield.

OpenStudy (michele_laino):

I think that the requested probability, is given by the subsequent ratio of areas: \[\Large p = \frac{{{\text{mound Area}}}}{{{\text{diamon Area}}}} \times 100\]

OpenStudy (anonymous):

Thank you.

OpenStudy (michele_laino):

oops.. diamond*

OpenStudy (anonymous):

lol

OpenStudy (michele_laino):

:)

OpenStudy (michele_laino):

\[\large p = \frac{{{\text{mound Area}}}}{{{\text{diamond Area}}}} \times 100 = \frac{{3.14 \times 9 \times 9}}{{90 \times 90}} \times 100 = ...\% \]

OpenStudy (anonymous):

Okay, I get it now.

OpenStudy (anonymous):

thank you sooo much

OpenStudy (michele_laino):

:)